OpenStudy (anonymous):

what is the y coordinate of the y−intercept of the line that passes through the points (−2, −2) and (2, 4)?

OpenStudy (anonymous):

 The answer is 2

OpenStudy (anonymous):

:D If ya need more help just tell me

OpenStudy (campbell_st):

find the equation of the line 1. the gradient m = rise/run (4 -(-2))/(2-(-2)) = = 6/4 or 3/2 using the point slope formula with the point (2, 4) y- 4 = 3/2(x - 2) y = 3/2x + 3 line is in the form y = mx + b b is the y-intercept the y intercept is 3